# Predicting Video Slot Attention Queries from Random Slot-Feature Pairs

### v2

@@ Introduction
By video Object-Centric Learning (OCL) (Singh, Deng, and Ahn 2022; Zadaianchuk, Seitzer, and Martius 2024), objects in the video can be discovered and represented as respective single feature vectors, termed slots, as well as tracked across frames. This is achieved under self-supervision, by simply forcing the slots to reconstruct the input video frames in some format. Representing a visual scene described by a video as object-level features is not only cognitively plausible as we humans perceive the world similarly (Palmeri and Gauthier 2004; Cavanagh 2011), but also practically feasible as a visual scene evolves upon dynamics among objects. This is why OCL is gaining popularity in applications like scene representing (Locatello et al. 2020), dynamics modeling (Wu et al. 2023a), planning and decision-making (Palmeri and Gauthier 2004; Ferraro et al. 2025) lately.
Mainstream video OCL methods (Singh, Wu, and Ahn 2022; Kipf et al. 2022) adopt a recurrent architecture, as shown in Figure 1 (a). Given some query vectors, the aggregator, a Slot Attention (Locatello et al. 2020) module, aggregates current video frame into slots, where each of such feature vectors represents an object. For next frame aggregation, the transitioner, typically a Transformer encoder block (Vaswani et al. 2017), transits current slots into the next query to provide temporally consistent representations for objects in dynamically evolving scenes, ensuring coherent tracking and identity preservation across time. Despite its overwhelmingly wide adoptation, does the query prediction via the transitioner between frames really learn the knowledge to effectively power OCL on videos?
@@
Figure 2: Our model architecture. (left) Our method is built upon SlotContrast (Manasyan et al. 2025). A frozen DINO2 (Oquab et al. 2023) model encodes current video frame $\boldsymbol{I}_{t}$ into current feature $\boldsymbol{F}_{t}$; A Slot Attention (Locatello et al. 2020) module aggregates $\boldsymbol{F}_{t}$ into current object-level vectors, slots $\boldsymbol{S}_{t}$, under current query $\boldsymbol{Q}_{t}$; A Transformer decoder block (Vaswani et al. 2017) transits $\boldsymbol{S}_{t}$ conditioned on next feature $\boldsymbol{F}_{t+1}$ to next query $\boldsymbol{Q}_{t+1}$; A random Transformer decoder (Zhao et al. 2025d) decodes $\boldsymbol{S}_{t}$ into current reconstruction $\boldsymbol{F}^{\prime}_{t}$. The objective is minimizing difference between $\boldsymbol{F}_{t}$ and $\boldsymbol{F}^{\prime}_{t}$. (right) How our transitioner works. To effectively learn transition dynamics, our transitioner explores slots $\boldsymbol{S}_{t_{1}}$ and feature $\boldsymbol{F}_{t_{2}}$ at any past time point within window size $\mathit{\Delta}$ to predict next query $\boldsymbol{Q}_{t+1}$ during training. Relative time embeddings are added to $\boldsymbol{S}_{t_{1}}$ and $\boldsymbol{F}_{t_{2}}$ to indicate their offset from $t+1$. To maximize prediction accuracy, our trainsitioner exploits only the most recent slots $\boldsymbol{S}_{t}$ and feature $\boldsymbol{F}_{t+1}$ to predict $\boldsymbol{Q}_{t+1}$ during evaluation. The left sub-figure is adapted from (Zhao et al. 2025c).
@@
Video object-centric learning. Given that Slot Attention has been central to video OCL since its inception, we classify existing methods as classical ones (Jiang et al. 2019; Kosiorek et al. 2018; van Steenkiste et al. 2018; Veerapaneni et al. 2020; Burgess et al. 2019) and modern ones (Kabra et al. 2021; Aydemir, Xie, and Guney 2023; Kipf et al. 2022; Elsayed et al. 2022; Safadoust and Güney 2023; Singh, Wu, and Ahn 2022; Traub et al. 2022; Zadaianchuk, Seitzer, and Martius 2024; Zoran et al. 2021; Qian et al. 2023). Modern methods’ performance has improved a lot even on real-world videos under self-supervision. Our work follows this line of research and we omit the word “modern” for simplicity. Video OCL is basically image OCL with recurrent module connecting between frames, i.e., current slots are transited to next query. Like image OCL, decoders, which are mixture-based (Locatello et al. 2020; Seitzer et al. 2023), auto-regressive (Singh, Deng, and Ahn 2022; Kakogeorgiou et al. 2024; Zhao et al. 2025d) or denoising-based (Jiang et al. 2023; Wu et al. 2023b), are also applicable to video OCL; Decoding targets, which are quantized with different inductive biases (Zhao et al. 2025a, b, c), are applicable to video OCL too. Specific to video OCL, temporal intrinsics like temporal prediction (Zadaianchuk, Seitzer, and Martius 2024) and consistency (Aydemir, Xie, and Guney 2023; Manasyan et al. 2025) can be utilized to enhance the performance. We focus on temporal inductive biases.
@@
Unsupervised video object tracking. Unsupervised Video Object Tracking (VOT) resembles (Li et al. 2019; Yuan et al. 2020; Lai, Lu, and Xie 2020; Wang et al. 2021; Xu and Wang 2021; Shen et al. 2022; Li et al. 2022) video OCL very much, because both are query-based and the query/slots should be maintained consistently across frames under self-supervision. However, VOT query only covers interested objects, while video OCL query covers all objects and the background. Essentially, unsupervised VOT only cares about temporal consistency of query representations while video OCL also demand slot representations to be informative enough to represent the scene. Video OCL slots, the sub-symbolic representations, can be directly employed in downstream scene representation and dynamics modeling, while VOT representations can not. Comparison with unsupervised VOT methods is of out scope.
@@
Query prediction. It is query prediction that adapts image OCL to video OCL (Singh, Wu, and Ahn 2022). Most video OCL methods employ a single Transformer encoder block to transit current slots into next query, i.e., the naive recurrent architecture of aggreagtion-transition-aggregation… (Zhao et al. 2025c). Well-recognized works like STEVE (Singh, Wu, and Ahn 2022), SAVi (Kipf et al. 2022), SAVi++ (Elsayed et al. 2022), VideoSAUR (Zadaianchuk, Seitzer, and Martius 2024) and SlotContrast (Manasyan et al. 2025) all take such design. There is no transitioner in SOLV (Aydemir, Xie, and Guney 2023) because global slots are aggregated for all frames, which however requires all frames to be available once together. A few works STATM (Li et al. 2025b) and SlotPi (Li et al. 2025a) fuse architectures of video OCL (Kipf et al. 2022) and world model (Wu et al. 2023a) and predict next query as auto-regression, which can be taken as an advanced recurrent architecture. However, none of them acknowledge or explore issues (i1) and (i2), which are directly addressed by our method.
@@
As shown in Figure 2 (left), we build our whole method upon SlotContrast (Manasyan et al. 2025), the latest state-of-the-art. Our model consists of an encoder $\boldsymbol{\phi}_{\mathrm{e}}$, an aggregator $\boldsymbol{\phi}_{\mathrm{a}}$, a decoder $\boldsymbol{\phi}_{\mathrm{d}}$ and a transitioner $\boldsymbol{\phi}_{\mathrm{t}}$.
@@
where $\boldsymbol{\phi}_{\mathrm{e}}$ is parameterized as a DINO2 (Oquab et al. 2023) model, which is a pretrained vision foundation model and is always frozen; $h_{0}$, $w_{0}$ and $c_{0}$ denote the input resolution, i.e., height, width and channel, while $h$, $w$ and $c$ are that after encoding.
@@
where $\boldsymbol{\phi}_{\mathrm{a}}$ is parameterized as typical Slot Attention (Locatello et al. 2020); $s$ is the number of query vectors; $\boldsymbol{M}_{t}$ is binarized from $\boldsymbol{S}_{t}$’s attention maps along the slot dimension. Because Slot Attention is a kind of attention, $\boldsymbol{Q}_{t}$ and $\boldsymbol{S}_{t}$ always have identical tensor shapes. $\boldsymbol{M}_{t}$ can be used for object discovery, while $\boldsymbol{S}_{t}$ can be used for downstream scene representation and dynamics modeling.
@@
where $\boldsymbol{\phi}_{\mathrm{r}}$ is parameterized as a single Transformer decoder block (Vaswani et al. 2017). Aggregator $\boldsymbol{\phi}_{\mathrm{a}}$ and transitioner $\boldsymbol{\phi}_{\mathrm{r}}$ form the aggregation-transition recurrence, which powers OCL through video frames. Please note that extra specific processes on $\boldsymbol{S}_{t}$ and $\boldsymbol{F}_{t+1}$ here are needed during training and evaluation, but are omitted for simplicity. We will detail them in the following two subsections.
@@
where $\boldsymbol{\phi}_{\mathrm{d}}$ is parameterized as a random auto-regressive Transformer decoder (Zhao et al. 2025d). The input clue needed in such decoding is omitted for simplicity.
@@
where $\mathrm{MSE}(\cdot,\cdot)$ is mean squared error; $\mathrm{sg}(\cdot)$ is stopping gradient; $T$ is the total number frames in a video. The auxiliary loss (Manasyan et al. 2025) is omitted for simplicity.
@@
where $\boldsymbol{\phi}_{\mathrm{r}}^{1}$ is parameterized as a Transformer encoder block (Vaswani et al. 2017; Singh, Wu, and Ahn 2022; Kipf et al. 2022). The transitioner $\boldsymbol{\phi}_{\mathrm{r}}^{2}$ of remaining video OCL methods predict from all available slots $\{\boldsymbol{S}_{i}\}_{i=0}^{t}$:
@@
where $\boldsymbol{\phi}_{\mathrm{r}}^{2}$ is parameterized as some multi-layer Transformer encoder variant (Li et al. 2025b, a).
@@
Here $\boldsymbol{E}\in\mathbb{R}^{\mathit{\Delta}\times c}$ is the same variable in Equation 6. We empirically set $\mathit{\Delta}$ to be identical to the widely-adopted video clip length in training, e.g., $\mathit{\Delta}=6$ given video length $T=24$ (Kipf et al. 2022; Elsayed et al. 2022) and $\mathit{\Delta}=5$ given video length $T=20$ (Zhao et al. 2025d).
@@
Baselines. STEVE (Singh, Wu, and Ahn 2022) realizes the first Slot Attention-based OCL on videos. VideoSAUR (Zadaianchuk, Seitzer, and Martius 2024) is the first method that utilizes vision foundation model and achieves competitive results in OCL on real-world complex videos. SlotContrast (Manasyan et al. 2025) enables temporal consistency on long videos, setting new state-or-the-art (SotA) recently. DIAS (Zhao et al. 2025d), another state-of-the-art, supports OCL on both images and videos and we use the latter version without its slot pruning trick for fair comparison. Our method RandSF.Q is compared with these methods. Comparing with methods like SOLV (Aydemir, Xie, and Guney 2023), which uses slot pruning, is unfair. So does comparing with SAVi (Kipf et al. 2022) and SAVi++ (Elsayed et al. 2022), which are trained with external stronger supervision like optical flow and depth map.
@@
roughly for background segmentation accuracy measurement, ARI foreground (ARIfg) for foreground large objects, mean Best Overlap (mBO) (Uijlings et al. 2013) for best overlapped regions, and mean Intersection over Union (mIoU) [^6] https://scikit-learn.org/stable/modules/generated/sklearn.metri
@@
Object recognition. We compare our method RandSF.Q with baseline SlotContrast in terms of their object recognition performance on dataset YTVIS. This consists of two sub-tasks, i.e., supervised classification and regression from slots. Some literature (Locatello et al. 2020) names this as set prediction. We follow (Seitzer et al. 2023) to represent the dataset as slots and train a two-layer MLP to predict the object class and bounding box corresponding to each slot, supervised by the annotations of object class labels and bounding boxes in the dataset. We use top1 accuracy [^7] https://scikit-learn.org/stable/modules/generated/sklearn.metri
@@
Visual question answering. We compare our method RandSF.Q with baseline SlotContrast in terms of their visual question answering performance on dataset CLEVRER. We follow the data processing and model design of Aloe (Ding et al. 2021; Wu et al. 2023a). We firstly pre-train OCL models on CLEVRER and then freeze them. Each video’s frames are all represented as slots, then slots of different frames are added with corresponding time step embeddings; meanwhile, question texts are embedded into vectors and added with corresponding position embeddings. These tokens are fed into Aloe together, appended with a classification token. The output is obtained by projecting the transformed classification token into logits of all answer labels.
@@
Object dynamics prediction. We focus on visual prediction as it has much less downstream factors that might confound with the results compared with other tasks. We compare our method RandSF.Q with baseline SlotContrast in term of their object dynamics prediction performance on dataset YTVIS. We follow the model architecturing settings from SlotFormer (Wu et al. 2023a). We first pre-train OCL models on YTVIS and freez them. Then we use the OCL model for representing video sample in YTVIS as slots, and input them into world model SlotFormer for predicting slots of future time steps. Like both (Wu et al. 2023a) and (Manasyan et al. 2025), we do not measure the reconstructed pixel errors from predicted slots; instead, we measure the byproduct segmentation accuracy accumulated to current frame from predicted slots. The former may demonstrate good reconstruction performance but poor object dynamics. In contrast, the latter directly reflects how well each object being predicted and represented.

### v4

@@ Introduction
By video Object-Centric Learning (OCL) (singh2021slate; zadaianchuk2024videosaur), objects in the video can be discovered and represented as respective single feature vectors, termed slots, as well as tracked across frames. This is achieved under self-supervision, by simply forcing the slots to reconstruct the input video frames in some format. Representing a visual scene described by a video as object-level features is not only cognitively plausible as we humans perceive the world similarly (palmeri2004visual; cavanagh2011visual), but also practically feasible as a visual scene evolves upon dynamics among objects. This is why OCL is gaining popularity in applications like scene representing (locatello2020slotattent), dynamics modeling (wu2022slotformer), planning and decision-making (palmeri2004visual; ferraro2025focus) lately.
Mainstream video OCL methods (singh2022steve; kipf2021savi) adopt a recurrent architecture, as shown in Figure 1 (a). Given some query vectors, the aggregator, a Slot Attention (locatello2020slotattent) module, aggregates current video frame into slots, where each of such feature vectors represents an object. For next frame aggregation, the transitioner, typically a Transformer encoder block (vaswani2017transformer), transits current slots into the next query to provide temporally consistent representations for objects in dynamically evolving scenes, ensuring coherent tracking and identity preservation across time. Despite its overwhelmingly wide adoptation, does the query prediction via the transitioner between frames really learn the knowledge to effectively power OCL on videos?
@@
Figure 2: Our model architecture. (left) Our method is built upon SlotContrast (manasyan2025slotcontrast). A frozen DINO2 (oquab2023dino2) model encodes current video frame $\boldsymbol{I}_{t}$ into current feature $\boldsymbol{F}_{t}$; A Slot Attention (locatello2020slotattent) module aggregates $\boldsymbol{F}_{t}$ into current object-level vectors, slots $\boldsymbol{S}_{t}$, under current query $\boldsymbol{Q}_{t}$; A Transformer decoder block (vaswani2017transformer) transits $\boldsymbol{S}_{t}$ conditioned on next feature $\boldsymbol{F}_{t+1}$ to next query $\boldsymbol{Q}_{t+1}$; A random Transformer decoder (zhao2025dias) decodes $\boldsymbol{S}_{t}$ into current reconstruction $\boldsymbol{F}^{\prime}_{t}$. The objective is minimizing difference between $\boldsymbol{F}_{t}$ and $\boldsymbol{F}^{\prime}_{t}$. (right) How our transitioner works. To effectively learn transition dynamics, our transitioner explores slots $\boldsymbol{S}_{t_{1}}$ and feature $\boldsymbol{F}_{t_{2}}$ at any past time point within window size $\mathit{\Delta}$ to predict next query $\boldsymbol{Q}_{t+1}$ during training. Relative time embeddings are added to $\boldsymbol{S}_{t_{1}}$ and $\boldsymbol{F}_{t_{2}}$ to indicate their offset from $t+1$. To maximize prediction accuracy, our trainsitioner exploits only the most recent slots $\boldsymbol{S}_{t}$ and feature $\boldsymbol{F}_{t+1}$ to predict $\boldsymbol{Q}_{t+1}$ during evaluation. The left sub-figure is adapted from (zhao2025vvo).
@@
Video object-centric learning. Given that Slot Attention has been central to video OCL since its inception, we classify existing methods as classical ones (jiang2019scalor; kosiorek2018sqair; van2018relationalnem; veerapaneni2020op3; burgess2019monet) and modern ones (kabra2021simone; aydemir2023solv; kipf2021savi; elsayed2022savipp; safadoust2023multi; singh2022steve; traub2022loci; zadaianchuk2024videosaur; zoran2021parts; qian2023smtc). Modern methods’ performance has improved a lot even on real-world videos under self-supervision. Our work follows this line of research and we omit the word “modern” for simplicity. Video OCL is basically image OCL with recurrent module connecting between frames, i.e., current slots are transited to next query. Like image OCL, decoders, which are mixture-based (locatello2020slotattent; seitzer2023dinosaur), auto-regressive (singh2021slate; kakogeorgiou2024spot; zhao2025dias) or denoising-based (jiang2023lsd; wu2023slotdiffuz), are also applicable to video OCL; Decoding targets, which are quantized with different inductive biases (zhao2025gdr; zhao2025msf; zhao2025vvo), are applicable to video OCL too. Specific to video OCL, temporal intrinsics like temporal prediction (zadaianchuk2024videosaur) and consistency (aydemir2023solv; manasyan2025slotcontrast) can be utilized to enhance the performance. We focus on temporal inductive biases.
@@
Unsupervised video object tracking. Unsupervised Video Object Tracking (VOT) resembles (li2019joint; yuan2020self; lai2020cvpr; wang2021unsupervised; xu2021iccv; shen2022cvpr; li2022cvpr) video OCL very much, because both are query-based and the query/slots should be maintained consistently across frames under self-supervision. However, VOT query only covers interested objects, while video OCL query covers all objects and the background. Essentially, unsupervised VOT only cares about temporal consistency of query representations while video OCL also demand slot representations to be informative enough to represent the scene. Video OCL slots, the sub-symbolic representations, can be directly employed in downstream scene representation and dynamics modeling, while VOT representations can not. Comparison with unsupervised VOT methods is of out scope.
@@
Query prediction. It is query prediction that adapts image OCL to video OCL (singh2022steve). Most video OCL methods employ a single Transformer encoder block to transit current slots into next query, i.e., the naive recurrent architecture of aggreagtion-transition-aggregation… (zhao2025vvo). Well-recognized works like STEVE (singh2022steve), SAVi (kipf2021savi), SAVi++ (elsayed2022savipp), VideoSAUR (zadaianchuk2024videosaur) and SlotContrast (manasyan2025slotcontrast) all take such design. There is no transitioner in SOLV (aydemir2023solv) because global slots are aggregated for all frames, which however requires all frames to be available once together. A few works STATM (li2025statm) and SlotPi (li2025slotpi) fuse architectures of video OCL (kipf2021savi) and world model (wu2022slotformer) and predict next query as auto-regression, which can be taken as an advanced recurrent architecture. However, none of them acknowledge or explore issues (i1) and (i2), which are directly addressed by our method.
@@
As shown in Figure 2 (left), we build our whole method upon SlotContrast (manasyan2025slotcontrast), the latest state-of-the-art. Our model consists of an encoder $\boldsymbol{\phi}_{\mathrm{e}}$, an aggregator $\boldsymbol{\phi}_{\mathrm{a}}$, a decoder $\boldsymbol{\phi}_{\mathrm{d}}$ and a transitioner $\boldsymbol{\phi}_{\mathrm{t}}$.
@@
where $\boldsymbol{\phi}_{\mathrm{e}}$ is parameterized as a DINO2 (oquab2023dino2) model, which is a pretrained vision foundation model and is always frozen; $h_{0}$, $w_{0}$ and $c_{0}$ denote the input resolution, i.e., height, width and channel, while $h$, $w$ and $c$ are that after encoding.
@@
where $\boldsymbol{\phi}_{\mathrm{a}}$ is parameterized as typical Slot Attention (locatello2020slotattent); $s$ is the number of query vectors; $\boldsymbol{M}_{t}$ is binarized from $\boldsymbol{S}_{t}$’s attention maps along the slot dimension. Because Slot Attention is a kind of attention, $\boldsymbol{Q}_{t}$ and $\boldsymbol{S}_{t}$ always have identical tensor shapes. $\boldsymbol{M}_{t}$ can be used for object discovery, while $\boldsymbol{S}_{t}$ can be used for downstream scene representation and dynamics modeling.
@@
where $\boldsymbol{\phi}_{\mathrm{r}}$ is parameterized as a single Transformer decoder block (vaswani2017transformer). Aggregator $\boldsymbol{\phi}_{\mathrm{a}}$ and transitioner $\boldsymbol{\phi}_{\mathrm{r}}$ form the aggregation-transition recurrence, which powers OCL through video frames. Please note that extra specific processes on $\boldsymbol{S}_{t}$ and $\boldsymbol{F}_{t+1}$ here are needed during training and evaluation, but are omitted for simplicity. We will detail them in the following two subsections.
@@
where $\boldsymbol{\phi}_{\mathrm{d}}$ is parameterized as a random auto-regressive Transformer decoder (zhao2025dias). The input clue needed in such decoding is omitted for simplicity.
@@
where $\mathrm{MSE}(\cdot,\cdot)$ is mean squared error; $\mathrm{sg}(\cdot)$ is stopping gradient; $T$ is the total number frames in a video. The auxiliary loss (manasyan2025slotcontrast) is omitted for simplicity.
@@
where $\boldsymbol{\phi}_{\mathrm{r}}^{1}$ is parameterized as a Transformer encoder block (vaswani2017transformer; singh2022steve; kipf2021savi). The transitioner $\boldsymbol{\phi}_{\mathrm{r}}^{2}$ of remaining video OCL methods predict from all available slots $\{\boldsymbol{S}_{i}\}_{i=0}^{t}$:
@@
where $\boldsymbol{\phi}_{\mathrm{r}}^{2}$ is parameterized as some multi-layer Transformer encoder variant (li2025statm; li2025slotpi).
@@
Here $\boldsymbol{E}\in\mathbb{R}^{\mathit{\Delta}\times c}$ is the same variable in Equation 6. We empirically set $\mathit{\Delta}$ to be identical to the widely-adopted video clip length in training, e.g., $\mathit{\Delta}=6$ given video length $T=24$ (kipf2021savi; elsayed2022savipp) and $\mathit{\Delta}=5$ given video length $T=20$ (zhao2025dias).
@@
Baselines. STEVE (singh2022steve) realizes the first Slot Attention-based OCL on videos. VideoSAUR (zadaianchuk2024videosaur) is the first method that utilizes vision foundation model and achieves competitive results in OCL on real-world complex videos. SlotContrast (manasyan2025slotcontrast) enables temporal consistency on long videos, setting new state-or-the-art (SotA) recently. DIAS (zhao2025dias), another state-of-the-art, supports OCL on both images and videos and we use the latter version without its slot pruning trick for fair comparison. Our method RandSF.Q is compared with these methods. Comparing with methods like SOLV (aydemir2023solv), which uses slot pruning, is unfair. So does comparing with SAVi (kipf2021savi) and SAVi++ (elsayed2022savipp), which are trained with external stronger supervision like optical flow and depth map.
@@
roughly for background segmentation accuracy measurement, ARI foreground (ARIfg) for foreground large objects, mean Best Overlap (mBO) (uijlings2013selectivesearch) for best overlapped regions, and mean Intersection over Union (mIoU) [^6] https://scikit-learn.org/stable/modules/generated/sklearn.metri
@@
Object recognition. We compare our method RandSF.Q with baseline SlotContrast in terms of their object recognition performance on dataset YTVIS. This consists of two sub-tasks, i.e., supervised classification and regression from slots. Some literature (locatello2020slotattent) names this as set prediction. We follow (seitzer2023dinosaur) to represent the dataset as slots and train a two-layer MLP to predict the object class and bounding box corresponding to each slot, supervised by the annotations of object class labels and bounding boxes in the dataset. We use top1 accuracy [^7] https://scikit-learn.org/stable/modules/generated/sklearn.metri
@@
Visual question answering. We compare our method RandSF.Q with baseline SlotContrast in terms of their visual question answering performance on dataset CLEVRER. We follow the data processing and model design of Aloe (ding2021aloe; wu2022slotformer). We firstly pre-train OCL models on CLEVRER and then freeze them. Each video’s frames are all represented as slots, then slots of different frames are added with corresponding time step embeddings; meanwhile, question texts are embedded into vectors and added with corresponding position embeddings. These tokens are fed into Aloe together, appended with a classification token. The output is obtained by projecting the transformed classification token into logits of all answer labels.
@@
Object dynamics prediction. We focus on visual prediction as it has much less downstream factors that might confound with the results compared with other tasks. We compare our method RandSF.Q with baseline SlotContrast in term of their object dynamics prediction performance on dataset YTVIS. We follow the model architecturing settings from SlotFormer (wu2022slotformer). We first pre-train OCL models on YTVIS and freez them. Then we use the OCL model for representing video sample in YTVIS as slots, and input them into world model SlotFormer for predicting slots of future time steps. Like both (wu2022slotformer) and (manasyan2025slotcontrast), we do not measure the reconstructed pixel errors from predicted slots; instead, we measure the byproduct segmentation accuracy accumulated to current frame from predicted slots. The former may demonstrate good reconstruction performance but poor object dynamics. In contrast, the latter directly reflects how well each object being predicted and represented.